OpenStudy (anonymous):

Is this correct? A six-sided die of unknown bias is rolled 20 times, and the number 3 comes up 6 times. In the next three rounds (the die is rolled 20 times in each round), the number 3 comes up 6 times, 5 times, and 7 times. The experimental probability of rolling a 3 is ? %, which is approximately ? % more than its theoretical probability. Round off your answers The probability of rolling a 3 is 30 % which is approximately 30 % more than its theoretical probability.

OpenStudy (anonymous):

18/60=30

OpenStudy (anonymous):

@Michele_Laino can you help me please

OpenStudy (michele_laino):

I think that the experimental probability, is given by the subsequent formula: \[probability = \frac{{favorable\;cases}}{{possible\;cases}} = \frac{{6 + 5 + 7}}{{10 + 10 + 10}} = ...?\]

OpenStudy (michele_laino):

oops... \[probability = \frac{{favorable\;cases}}{{possible\;cases}} = \frac{{6 + 5 + 7}}{{20 + 20 + 20}} = ...?\]

OpenStudy (michele_laino):

more precisely, since we have rolled our die four rounds, then the right formula is: \[probability = \frac{{favorable\;cases}}{{possible\;cases}} = \frac{{6 + 6 + 5 + 7}}{{20 + 20 + 20 + 20}} = ...?\]

OpenStudy (anonymous):

9/40

OpenStudy (anonymous):

22.5%

OpenStudy (michele_laino):

I think 24/80=30%

OpenStudy (anonymous):

so 4 rounds because they say in the next 3. I missed that

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

our experiment includes four rounds

OpenStudy (anonymous):

ok I see but where did you het the extra 6 from sorry confused

OpenStudy (michele_laino):

first 6 is from the first round

OpenStudy (anonymous):

nevermind I reread it. Im terrible at these

OpenStudy (anonymous):

Thanks so much!!!!

OpenStudy (michele_laino):

Thanks!!!